OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x - 4)(2x - 1)(x + 7)?

OpenStudy (anonymous):

x={4,1/2,-7}

OpenStudy (anonymous):

(x-4)=0 x=4 (2x-1)=0 x=1/2 (x+7)=0 x=-7